How many consecutive integers beginning with 5 must be taken for the sum to equal 95.

Answer by FrankM(1040)

The series has N terms. The first term is 5, the last term is 5+(N-1) which simplifies to N+4
The sum of a series of consecutive integers is N/2 (first term +last term) so we have


(N)(5+(N+4))/2=95
(N)(N+9)=190
N^2+9N-190=0
(N-10)(N+19)=0
N=10
10 terms 5 thru 14 add to 95